# Approximating Unrelated Machine Weighted Completion Time Using Iterative Rounding and Computer Assisted Proofs

### v1

@@ 2.2 Configuration LP
$\sum_{f}y_{if}$ $=1$ $\forall i\in M$ (2) $\sum_{i\in M}\sum_{f\ni j}y_{if}$ $=1$ $\forall j\in J$ (3) $y_{if}$ $\geq 0$ $\forall i\in M,j\in J$ (4)
(2) requires that for every machine $i\in M$, we choose exactly one configuration $f$. (3) requires that every job is covered by exactly one configuration across all machines. (4) is the non-negativity constraint. We pay a cost of $\mathrm{cost}_{i}(f)$ on $i$ if the set of jobs assigned to $i$ is $f$, thus we have the objective (1). Though this LP has exponential number of variables, Sviridenko and Wiese [23] showed that it can be solved approximately within a factor of $1+\epsilon$, for any constant $\epsilon>0$. So, we assume we are given a $(1+\epsilon)$-approximate solution $y$ to the LP; in particular, the result of [23] suggests that the number of non-zeros in $y$ is small. For every $i\in M$ and $j\in J$, we let $z_{ij}:=\sum_{f\ni j}y_{if}$ denote the fraction of job $j$ assigned to machine $i$. We only use $z$ variables in the rounding algorithm, while reserving the $y_{if}$ variables for the analysis.
@@ 2.3 Partition of Jobs into Classes and Construction of Bipartite Graph $G$ with Marked and Unmarked Edges
Let $\rho$ be a constant which will be set to $2$ later; for future reference we present the algorithm and a part of the analysis for a general $\rho>1$. We randomly choose a $\beta\in[1,\rho)$ such that $\ln\beta$ is uniformly distributed in $[0,\ln\rho)$. For an integer $k$, we say a job $j$ is in class $k$ if $p_{j}\in[\beta\rho^{k},\beta\rho^{k+1})$. Let $J_{k}$ be the set of jobs in class $k$; notice that it depends on the randomly chosen $\beta$. Once $\beta$ is chosen, we have a global partition $(J_{k})_{k}$ of jobs, independent of the machines. We construct a bipartite multi-graph $G=(M,J,E)$ between $M$ and $J$, along with a vector $x\in(0,1]^{E}$. Each edge in $E$ is either marked or unmarked. The procedure is formally described in Algorithm 1.
@@
10:         elsenormal-$\triangleright$ We have $v<\beta\rho^{k}<v+p_{j}z_{ij}$
@@
Notations
With $G=(M,J,E)$ and $x\in(0,1]^{E}$ constructed, we define the following notations. For every $e=ij\in E$, we define $p_{e}=p_{j}$ to be the size of its incident job, and $\sigma_{e}=\frac{w_{ij}}{p_{j}}$ be the Smith ratio. For every $i\in M$ (resp., $j\in J$), let $\delta_{i}$ (resp., $\delta_{j}$) be the set of edges in $E$ incident to $i$ (resp., $j$). For every job class $k$, let $E_{k}$ be the set of edges in $E$ between $M$ and $J_{k}$, $G_{k}=(M,J_{k},E_{k})$ and $x^{(k)}\in(0,1]^{E_{k}}$ be the $x$ vector restricted to $E_{k}$. For every $k$ and $i\in M$, let $\delta^{k}_{i}=\delta_{i}\cap E_{k}$ be incident edges of $i$ in $E_{k}$. We use superscripts “$\mathrm{mk}$” and “$\mathrm{umk}$” to denote the restrictions to marked and unmarked edges respectively: $E^{\mathrm{mk}}$ and $E^{\mathrm{umk}}$ are the sets of marked and unmarked edges in $E$ respectively. $\delta^{\mathrm{mk}}_{i}:=E^{\mathrm{mk}}\cap\delta_{i},\delta^{k,\mathrm{mk}}_{i}:=E^{\mathrm{mk}}\cap\delta^{k}_{i}$ and $\delta^{\mathrm{mk}}_{j}=E^{\mathrm{mk}}\cap\delta_{j}$; the same rule applies to unmarked edges as well.
@@ 2.4 Iterative Rounding
In words, the support of $a$ is structure. For a job $j$ on the structure, the sum of $a$-values of its two incident edges on the cycle is 00. If the structure is a cycle, then for any machine $i$ on the cycle, the sum of $a_{ij}p_{j}$ over its two incident edges $ij$ on the cycle is 00. If the structure is a pseudo-marked-path, we require the equality to hold for machines in $\{i_{1},i_{2},\cdots,i_{t-1}\}$ and its two incident marked edges on the path. It is easy to see that such $a$ exists and is unique up to scaling. The properties for the $a$ vector constructed in either case are summarized in the following claim.
@@ 2.5 Helper Lemmas
If the structure we found in the iteration is a cycle, then $a_{e}=0$ as $e$ is unmarked. The lemma follows from Claim 2.6. So, we assume the structure we found is a pseudo-marked-path. By Lemma 2.9, if at the beginning of the iteration, $|\delta^{k,\mathrm{mk}}_{i}\cap\bar{E}|\geq 2$, then $\sum_{e^{\prime}\in\delta^{k,\mathrm{mk}}_{i}}\bar{x}^{\mathrm{new}}_{e^{\prime}}p_{e^{\prime}}=\sum_{e^{\prime}\in\delta^{k,\mathrm{mk}}_{i}}\bar{x}^{\mathrm{old}}_{e^{\prime}}p_{e^{\prime}}$ always holds.
@@ 2.6 Algorithm 2 Returns an Integral Assignment
When the algorithm terminates, we do not have a cycle of marked edges, or a pseudo-marked-path in $\bar{E}$. So $(M,J_{k},E^{\mathrm{mk}}\cap\bar{E})$ is a forest. We focus on a non-singleton tree $T$ of marked edges in the forest. Let $J_{T}$ be the set of jobs in the tree $T$, and $M_{T}$ be the set of machines. Notice that $J_{T}$ is incident to at most 1 unmarked edge in $\bar{E}$: If there are two unmarked edges $ji$ and $j^{\prime}i^{\prime}$ in $\bar{E}$ with $j,j^{\prime}\in J_{T}$, then we can take the path from $j$ to $j^{\prime}$ in $T$, and concatenate it with $ij$ at the beginning and with $j^{\prime}i^{\prime}$ at the end. This would give us a pseudo-marked-path. First consider the case where there are no unmarked edges incident to $J_{T}$ in $\bar{G}$. Notice that $T$ contains at least 2 leaves. If it contains 2 leaf machines, then the path in $T$ between the two machines would be a pseudo-marked-path. Otherwise, $T$ contains at least one leaf-job $j\in J_{k}$. Assume $j$ is incident to $i\in M$ in $T$. If $i$ has degree at least $2$ in $T$, then $\sum_{e\in\delta^{k,\mathrm{mk}}_{i}}\bar{x}_{e}p_{e}=\sum_{e\in\delta^{k,\mathrm{mk}}_{i}}x_{e}p_{e}=\mathrm{vol}(\delta^{k,\mathrm{mk}}_{i})\leq\beta\rho^{k}$ by Property (2.11d) and Claim 2.3. Notice that $p_{j}\geq\beta\rho^{k}$ and $ij$ is the only edge incident to $j$ in $\bar{E}$, which implies $\bar{x}_{ij}=1$. This contradicts that $i$ has degree at least 2 in $T$. So $i$ has degree 1 in $T$ and $T$ consists of the single edge $ij$ with $x_{ij}=1$. Consider the second case where there is exactly one unmarked edge incident to $J_{T}$ in $\bar{G}$. Assume the edge is $e$ and it is incident to $j\in J_{T}$. If there is one leaf machine in $T$, then concatenating the path between the leaf machine and $j$ in $T$ and $e$ would give us a pseudo-marked-path. So $T$ does not have a leaf machine, and this implies $|J_{T}|\geq|M_{T}|+1$. Then again by Property (2.11d) and Claim 2.3 we have $\sum_{e\in\delta^{k,\mathrm{mk}}_{i}}\bar{x}_{e}p_{e}\leq\beta\rho^{k}$ for every machine $i\in M_{T}$. Every job in $J_{T}$ has size at least $\beta\rho^{k}$, and $J_{T}$ is incident to only 1 unmarked edge. This can only happen if $T$ is the singleton ($(\{j\},\emptyset)$. So, all the marked edges $e\in E_{k}$ have $\bar{x}_{e}\in\{0,1\}$. As every $j\in J_{k}$ can be incident to at most 1 unmarked edge in $\bar{E}$ in the end, all the unmarked edges $e\in E^{k}$ also have $\bar{x}_{e}\in\{0,1\}$. ∎
@@ 3 An Upper Bound on Weighted Completion Time on Machine $i$
In this section we fix a machine $i$, rewrite the LP cost on $i$, and prove an upper bound on the expected weighted completion time on $i$ in our solution. Then in Section 4, we analyze the ratio between the two quantities. As the edges $E$ and the vector $x\in(0,1]^{E}$ depend on the random choice $\beta$, in Section 4 it will be more convenient to express the formulations using jobs $J$ and the $z$-vector. So, we define $\sigma_{j}:=\frac{w_{ij}}{p_{j}}$ to be the Smith ratio of $j$ on machine $i$ for every $j\in J$. We index $j$ as $[n]$ so that $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{n}$. Define $\sigma_{n+1}=0$ for convenience. We first rewrite the LP cost on $i$:
@@
The last inequality used $\sum_{f\ni j}y_{if}=z_{ij}$ for every $j\in[n]$. ∎
Till the end of the section, we fix the random choice $\beta\in[1,\rho)$. Let ${\mathrm{wC}}^{(i)}$ denote total weighted completion time of jobs assigned to $i$, in the solution produced by our algorithm. Recall that $\widehat{\operatorname*{\mathbb{E}}}[\cdot]$ is $\operatorname*{\mathbb{E}}[\cdot|\beta]$. For a job $j\in[n]$, define $\mathrm{vol}(j):=z_{ij}p_{j}$. This is equal to $\sum_{e=ij}\mathrm{vol}(e)$. For a set $J^{\prime}\subseteq[n]$ of jobs, define $\mathrm{vol}(J^{\prime})=\sum_{j\in J^{\prime}}\mathrm{vol}(j)$. The main goal of this section is to prove the following lemma, that bounds the cost on $i$ in our solution:
@@
In the proof, we use edges and $x$-vector instead of jobs and $z$-vector. Recall that $\sigma_{e}=\frac{w_{ij}}{p_{j}}$ if $e$ is an edge between $i$ and $j$. We define a total order $\prec$ over the set $\delta_{i}$ of incident of edges of $i$ in $G$ using their order of creation in Algorithm 1: $e\prec e^{\prime}$ if $e$ is created before $e^{\prime}$; this implies $\sigma_{e}\geq\sigma_{e^{\prime}}$. In particular, for any two parallel edges $e$ and $e^{\prime}$ with $e$ marked and $e^{\prime}$ unmarked, we have $e\prec e^{\prime}$. We define $e\preceq e^{\prime}$ if $e\prec e^{\prime}$ or $e=e^{\prime}$. For every $e\in\delta^{i}$, we define $\mathrm{next}(e)$ to be the edge after $e$ in the order $\prec$. For the last edge $e$ in the order, we define $\sigma_{\mathrm{next}(e)}=0$ for convenience. We use $[e]:=\{e^{\prime}\in\delta_{i}:e^{\prime}\preceq e\}$ to denote the set of edges before $e$ in the order $\prec$, including the edge $e$ itself. With the notations defined, (6) becomes
$\widehat{\operatorname*{\mathbb{E}}}\big{[}{\mathrm{wC}}^{(i)}\big{]}\leq\sum_{e^{*}\in\delta_{i}}(\sigma_{e^{*}}-\sigma_{\mathrm{next}(e^{*})})\Big{(}\sum_{e\preceq e^{*}}x_{e}p_{e}^{2}+\frac{1}{2}\cdot\mathrm{vol}^{2}([e^{*}])-\frac{1}{2}\sum_{k}\big{(}\min\big{\{}\mathrm{vol}([e^{*}]\cap\delta^{i}_{k}),\beta\rho^{k}\big{\}}\big{)}^{2}\Big{)}.$ (7)
To see the equivalence of the two inequalities, we make two modifications to the right side (7). First, merging two parallel edges $e$ and $e^{\prime}$ into a new edge $e^{\prime\prime}$ with $x_{e^{\prime\prime}}=x_{e}+x_{e^{\prime}}$ does not change the quantity. ($e$ and $e^{\prime}$ are next to each other in the order $\prec$, and thus the new order can be naturally defined.) This holds as $\sigma_{e}=\sigma_{e^{\prime}},p_{e}=p_{e^{\prime}}$. Then for any $j^{*}\in[n]$ with no edges between $i$ and $j$ in $G$, we add an dummy edge $e=ij$ of $p_{e}=p_{J},\sigma_{e}=\frac{w_{ij}}{p_{j}},x_{e}=0$, and insert it into the order $\prec$ naturally. The operation does not change (7). After the two operations, the right side of (7) becomes (6). So now we focus on the proof of (7). For every $k$ and every $e\in\delta^{i}_{k}$, we define $X_{e}$ to be the final value of $\bar{x}_{e}$ in Algorithm 2 for the job class $k$. By Lemma 2.12, $X_{e}\in\{0,1\}$. The total weighted completion time of jobs on machine $i$ is
@@
The last equality used that $X_{e}\in\{0,1\}$, which implies $X_{e}^{2}=X_{e}$, for every $e\in\delta_{i}$. We fix an $e^{*}\in\delta_{i}$ and upper bound $\widehat{\operatorname*{\mathbb{E}}}\Big{[}\sum_{e^{\prime}\prec e\preceq e^{*}}X_{e}X_{e^{\prime}}p_{e}p_{e^{\prime}}\Big{]}$. Focus on two edges $e^{\prime}$ and $e$ with $e^{\prime}\prec e\preceq e^{*}$. First, if $e\in\delta^{k}_{i}$ and $e^{\prime}\in\delta^{k^{\prime}}_{i}$ for some $k\neq k^{\prime}$, then we have $\widehat{\operatorname*{\mathbb{E}}}[X_{e}X_{e^{\prime}}]=\widehat{\operatorname*{\mathbb{E}}}[X_{e}]\widehat{\operatorname*{\mathbb{E}}}[X_{e^{\prime}}]=x_{e}x_{e^{\prime}}$ as Algorithm 2 for job classes $k$ and $k^{\prime}$ are independent. So we assume $e^{\prime}$ and $e$ are both in $\delta^{k}_{i}$ for some $k$. If $e$ is marked, so is $e^{\prime}$. By Lemma 2.13, $X_{e}X_{e^{\prime}}=0$ always happens conditioned on the given $\beta$. If both $e$ and $e^{\prime}$ are unmarked, we have $\widehat{\operatorname*{\mathbb{E}}}[X_{e}X_{e^{\prime}}]\leq x_{e}x_{e^{\prime}}$ by Property (2.11c). For an edge $e\in[e^{*}]\cap\delta^{k,\mathrm{umk}}_{i}$, we have $\widehat{\operatorname*{\mathbb{E}}}\big{[}\sum_{e^{\prime}\in\delta^{k,\mathrm{mk}}_{i}}X_{e}X_{e^{\prime}}p_{e^{\prime}}\big{]}\leq\sum_{e^{\prime}\in\delta^{k,\mathrm{mk}}_{i}}x_{e}x_{e^{\prime}}p_{e^{\prime}}$ by Property (2.11e). Notice that $e^{\prime}\prec e$ for every $e^{\prime}\in\delta^{k,\mathrm{mk}}_{i}$. So, for an edge $e\in[e^{*}]\cap\delta^{k,\mathrm{umk}}_{i}$, we have
@@
The last equality follows from that we mark the first $\beta\rho^{k}$ volume of edges in $\sigma^{k}_{i}$ in Algorithm 1 according to the order $\prec$. Combining the inequality, (8) and that $\widehat{\operatorname*{\mathbb{E}}}\big{[}\sum_{e\preceq e^{*}}X_{e}p_{e}^{2}\big{]}=\sum_{e\preceq e^{*}}x_{e}p_{e}^{2}$, we have

### v2

@@ 2.2 Configuration LP
$\sum_{f}y_{if}$ $=1$ $\forall i\in M$ (2)
$\sum_{i\in M}\sum_{f\ni j}y_{if}$ $=1$ $\forall j\in J$ (3)
$y_{if}$ $\geq 0$ $\forall i\in M,j\in J$ (4)
(2) requires that for every machine $i\in M$, we choose exactly one configuration $f$. (3) requires that every job is covered by exactly one configuration across all machines. (4) is the non-negativity constraint. We pay a cost of $\mathrm{cost}_{i}(f)$ on $i$ if the set of jobs assigned to $i$ is $f$, thus we have the objective (1).
Though this LP has exponential number of variables, Sviridenko and Wiese [23] showed that it can be solved approximately within a factor of $1+\epsilon$, for any constant $\epsilon>0$. So, we assume we are given a $(1+\epsilon)$-approximate solution $y$ to the LP; in particular, the result of [23] suggests that the number of non-zeros in $y$ is small.
For every $i\in M$ and $j\in J$, we let $z_{ij}:=\sum_{f\ni j}y_{if}$ denote the fraction of job $j$ assigned to machine $i$. We only use $z$ variables in the rounding algorithm, while reserving the $y_{if}$ variables for the analysis.
Remark  We remark that the configuration LP could be replaced by the time-indexed LP relaxation in [17], where we have variables $x_{ijs}$ indicating if job $j$ is scheduled on machine $i$, with starting time $s$. A solution for the configuration LP can be derived from the time-index LP with one caveat: a configuration now is a multi-set of jobs. This does not pose a significant challenge to our algorithm and analysis, although we need to be careful in notations when handling multi-sets.
@@ 2.3 Partition of Jobs into Classes and Construction of Bipartite Graph $G$ with Marked and Unmarked Edges
Let $\rho$ be a constant which will be set to $2$ later; for future reference we present the algorithm and a part of the analysis for a general $\rho>1$. We randomly choose a $\beta\in[1,\rho)$ such that $\ln\beta$ is uniformly distributed in $[0,\ln\rho)$. For an integer $k$, we say a job $j$ is in class $k$ if $p_{j}\in[\beta\rho^{k},\beta\rho^{k+1})$. Let $J_{k}$ be the set of jobs in class $k$; notice that it depends on the randomly chosen $\beta$. Once $\beta$ is chosen, we have a global partition $(J_{k})_{k}$ of jobs, independent of the machines.
We construct a bipartite multi-graph $G=(M,J,E)$ between $M$ and $J$, along with a vector $x\in(0,1]^{E}$. Each edge in $E$ is either marked or unmarked. The procedure is formally described in Algorithm 1.
@@
10:         $\triangleright$ We have $v<\beta\rho^{k}<v+p_{j}z_{ij}$
@@
Notations  With $G=(M,J,E)$ and $x\in(0,1]^{E}$ constructed, we define the following notations. For every $e=ij\in E$, we define $p_{e}=p_{j}$ to be the size of its incident job, and $\sigma_{e}=\frac{w_{ij}}{p_{j}}$ be the Smith ratio. For every $i\in M$ (resp., $j\in J$), let $\delta_{i}$ (resp., $\delta_{j}$) be the set of edges in $E$ incident to $i$ (resp., $j$). For every job class $k$, let $E_{k}$ be the set of edges in $E$ between $M$ and $J_{k}$, $G_{k}=(M,J_{k},E_{k})$ and $x^{(k)}\in(0,1]^{E_{k}}$ be the $x$ vector restricted to $E_{k}$. For every $k$ and $i\in M$, let $\delta^{k}_{i}=\delta_{i}\cap E_{k}$ be incident edges of $i$ in $E_{k}$. We use superscripts “$\mathrm{mk}$” and “$\mathrm{umk}$” to denote the restrictions to marked and unmarked edges respectively: $E^{\mathrm{mk}}$ and $E^{\mathrm{umk}}$ are the sets of marked and unmarked edges in $E$ respectively. $\delta^{\mathrm{mk}}_{i}:=E^{\mathrm{mk}}\cap\delta_{i},\delta^{k,\mathrm{mk}}_{i}:=E^{\mathrm{mk}}\cap\delta^{k}_{i}$ and $\delta^{\mathrm{mk}}_{j}=E^{\mathrm{mk}}\cap\delta_{j}$; the same rule applies to unmarked edges as well.
@@ 2.4 Iterative Rounding
In words, the support of $a$ is structure. For a job $j$ on the structure, the sum of $a$-values of its two incident edges on the cycle is 00. If the structure is a cycle, then for any machine $i$ on the cycle, the sum of $a_{ij}p_{j}$ over its two incident edges $ij$ on the cycle is 00. If the structure is a pseudo-marked-path, we require the equality to hold for machines in $\{i_{1},i_{2},\cdots,i_{t-1}\}$ and its two incident marked edges on the path. It is easy to see that such $a$ exists and is unique up to scaling.
The properties for the $a$ vector constructed in either case are summarized in the following claim.
@@ 2.5 Helper Lemmas
If the structure we found in the iteration is a cycle, then $a_{e}=0$ as $e$ is unmarked. The lemma follows from Claim 2.6. So, we assume the structure we found is a pseudo-marked-path.
By Lemma 2.9, if at the beginning of the iteration, $|\delta^{k,\mathrm{mk}}_{i}\cap\bar{E}|\geq 2$, then $\sum_{e^{\prime}\in\delta^{k,\mathrm{mk}}_{i}}\bar{x}^{\mathrm{new}}_{e^{\prime}}p_{e^{\prime}}=\sum_{e^{\prime}\in\delta^{k,\mathrm{mk}}_{i}}\bar{x}^{\mathrm{old}}_{e^{\prime}}p_{e^{\prime}}$ always holds.
@@ 2.6 Algorithm 2 Returns an Integral Assignment
When the algorithm terminates, we do not have a cycle of marked edges, or a pseudo-marked-path in $\bar{E}$. So $(M,J_{k},E^{\mathrm{mk}}\cap\bar{E})$ is a forest. We focus on a non-singleton tree $T$ of marked edges in the forest. Let $J_{T}$ be the set of jobs in the tree $T$, and $M_{T}$ be the set of machines. Notice that $J_{T}$ is incident to at most 1 unmarked edge in $\bar{E}$: If there are two unmarked edges $ji$ and $j^{\prime}i^{\prime}$ in $\bar{E}$ with $j,j^{\prime}\in J_{T}$, then we can take the path from $j$ to $j^{\prime}$ in $T$, and concatenate it with $ij$ at the beginning and with $j^{\prime}i^{\prime}$ at the end. This would give us a pseudo-marked-path.
First consider the case where there are no unmarked edges incident to $J_{T}$ in $\bar{G}$. Notice that $T$ contains at least 2 leaves. If it contains 2 leaf machines, then the path in $T$ between the two machines would be a pseudo-marked-path. Otherwise, $T$ contains at least one leaf-job $j\in J_{k}$. Assume $j$ is incident to $i\in M$ in $T$. If $i$ has degree at least $2$ in $T$, then $\sum_{e\in\delta^{k,\mathrm{mk}}_{i}}\bar{x}_{e}p_{e}=\sum_{e\in\delta^{k,\mathrm{mk}}_{i}}x_{e}p_{e}=\mathrm{vol}(\delta^{k,\mathrm{mk}}_{i})\leq\beta\rho^{k}$ by Property (2.11d) and Claim 2.3. Notice that $p_{j}\geq\beta\rho^{k}$ and $ij$ is the only edge incident to $j$ in $\bar{E}$, which implies $\bar{x}_{ij}=1$. This contradicts that $i$ has degree at least 2 in $T$. So $i$ has degree 1 in $T$ and $T$ consists of the single edge $ij$ with $x_{ij}=1$.
Consider the second case where there is exactly one unmarked edge incident to $J_{T}$ in $\bar{G}$. Assume the edge is $e$ and it is incident to $j\in J_{T}$. If there is one leaf machine in $T$, then concatenating the path between the leaf machine and $j$ in $T$ and $e$ would give us a pseudo-marked-path. So $T$ does not have a leaf machine, and this implies $|J_{T}|\geq|M_{T}|+1$. Then again by Property (2.11d) and Claim 2.3 we have $\sum_{e\in\delta^{k,\mathrm{mk}}_{i}}\bar{x}_{e}p_{e}\leq\beta\rho^{k}$ for every machine $i\in M_{T}$. Every job in $J_{T}$ has size at least $\beta\rho^{k}$, and $J_{T}$ is incident to only 1 unmarked edge. This can only happen if $T$ is the singleton ($(\{j\},\emptyset)$.
So, all the marked edges $e\in E_{k}$ have $\bar{x}_{e}\in\{0,1\}$. As every $j\in J_{k}$ can be incident to at most 1 unmarked edge in $\bar{E}$ in the end, all the unmarked edges $e\in E^{k}$ also have $\bar{x}_{e}\in\{0,1\}$. ∎
@@ 3 An Upper Bound on Weighted Completion Time on Machine $i$
In this section we fix a machine $i$, rewrite the LP cost on $i$, and prove an upper bound on the expected weighted completion time on $i$ in our solution. Then in Section 4, we analyze the ratio between the two quantities. As the edges $E$ and the vector $x\in(0,1]^{E}$ depend on the random choice $\beta$, in Section 4 it will be more convenient to express the formulations using jobs $J$ and the $z$-vector. So, we define $\sigma_{j}:=\frac{w_{ij}}{p_{j}}$ to be the Smith ratio of $j$ on machine $i$ for every $j\in J$. We index $j$ as $[n]$ so that $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{n}$. Define $\sigma_{n+1}=0$ for convenience.
We first rewrite the LP cost on $i$:
@@
The last equality used $\sum_{f\ni j}y_{if}=z_{ij}$ for every $j\in[n]$. ∎
Till the end of the section, we fix the random choice $\beta\in[1,\rho)$. Let ${\mathrm{wC}}^{(i)}$ denote the total weighted completion time of jobs assigned to $i$ in the solution produced by our algorithm. Recall that $\widehat{\operatorname*{\mathbb{E}}}[\cdot]$ is $\operatorname*{\mathbb{E}}[\cdot|\beta]$. For a job $j\in[n]$, define $\mathrm{vol}(j):=z_{ij}p_{j}$. This is equal to $\sum_{e=ij}\mathrm{vol}(e)$. For a set $J^{\prime}\subseteq[n]$ of jobs, define $\mathrm{vol}(J^{\prime})=\sum_{j\in J^{\prime}}\mathrm{vol}(j)$. The main goal of this section is to prove the following lemma, that bounds the cost on $i$ in our solution:
@@
In the proof, we use edges and $x$-vector instead of jobs and $z$-vector. Recall that $\sigma_{e}=\frac{w_{ij}}{p_{j}}$ if $e$ is an edge between $i$ and $j$. We define a total order $\prec$ over the set $\delta_{i}$ of incident of edges of $i$ in $G$ using their order of creation in Algorithm 1: $e\prec e^{\prime}$ if $e$ is created before $e^{\prime}$; this implies $\sigma_{e}\geq\sigma_{e^{\prime}}$. In particular, for any two parallel edges $e$ and $e^{\prime}$ with $e$ marked and $e^{\prime}$ unmarked, we have $e\prec e^{\prime}$. We define $e\preceq e^{\prime}$ if $e\prec e^{\prime}$ or $e=e^{\prime}$. For every $e\in\delta_{i}$, we define $\mathrm{next}(e)$ to be the edge after $e$ in the order $\prec$. For the last edge $e$ in the order, we define $\sigma_{\mathrm{next}(e)}=0$ for convenience. We use $[e]:=\{e^{\prime}\in\delta_{i}:e^{\prime}\preceq e\}$ to denote the set of edges before $e$ in the order $\prec$, including the edge $e$ itself.
With the notations defined, (6) becomes
$\widehat{\operatorname*{\mathbb{E}}}\big{[}{\mathrm{wC}}^{(i)}\big{]}\leq\sum_{e^{*}\in\delta_{i}}(\sigma_{e^{*}}-\sigma_{\mathrm{next}(e^{*})})\Big{(}\sum_{e\preceq e^{*}}x_{e}p_{e}^{2}+\frac{1}{2}\cdot\mathrm{vol}^{2}([e^{*}])-\frac{1}{2}\sum_{k}\big{(}\min\big{\{}\mathrm{vol}([e^{*}]\cap\delta^{k}_{i}),\beta\rho^{k}\big{\}}\big{)}^{2}\Big{)}.$ (7)
To see the equivalence of the two inequalities, we make two modifications to the right side (7). First, merging two parallel edges $e$ and $e^{\prime}$ into a new edge $e^{\prime\prime}$ with $x_{e^{\prime\prime}}=x_{e}+x_{e^{\prime}}$ does not change the quantity. ($e$ and $e^{\prime}$ are next to each other in the order $\prec$, and thus the new order can be naturally defined.) This holds as $\sigma_{e}=\sigma_{e^{\prime}},p_{e}=p_{e^{\prime}}$. Then for any $j^{*}\in[n]$ with no edges between $i$ and $j$ in $G$, we add an dummy edge $e=ij$ of $p_{e}=p_{J},\sigma_{e}=\frac{w_{ij}}{p_{j}},x_{e}=0$, and insert it into the order $\prec$ naturally. The operation does not change (7). After the two operations, the right side of (7) becomes (6). So now we focus on the proof of (7).
For every $k$ and every $e\in\delta^{k}_{i}$, we define $X_{e}$ to be the final value of $\bar{x}_{e}$ in Algorithm 2 for the job class $k$. By Lemma 2.12, $X_{e}\in\{0,1\}$. The total weighted completion time of jobs on machine $i$ is
@@
The last equality used that $X_{e}\in\{0,1\}$, which implies $X_{e}^{2}=X_{e}$, for every $e\in\delta_{i}$. We fix an $e^{*}\in\delta_{i}$ and upper bound $\widehat{\operatorname*{\mathbb{E}}}\Big{[}\sum_{e^{\prime}\prec e\preceq e^{*}}X_{e}X_{e^{\prime}}p_{e}p_{e^{\prime}}\Big{]}$.
Focus on two edges $e^{\prime}$ and $e$ with $e^{\prime}\prec e\preceq e^{*}$. First, if $e\in\delta^{k}_{i}$ and $e^{\prime}\in\delta^{k^{\prime}}_{i}$ for some $k\neq k^{\prime}$, then we have $\widehat{\operatorname*{\mathbb{E}}}[X_{e}X_{e^{\prime}}]=\widehat{\operatorname*{\mathbb{E}}}[X_{e}]\widehat{\operatorname*{\mathbb{E}}}[X_{e^{\prime}}]=x_{e}x_{e^{\prime}}$ as Algorithm 2 for job classes $k$ and $k^{\prime}$ are independent. So we assume $e^{\prime}$ and $e$ are both in $\delta^{k}_{i}$ for some $k$. If $e$ is marked, so is $e^{\prime}$. By Lemma 2.13, $X_{e}X_{e^{\prime}}=0$ always happens conditioned on the given $\beta$. If both $e$ and $e^{\prime}$ are unmarked, we have $\widehat{\operatorname*{\mathbb{E}}}[X_{e}X_{e^{\prime}}]\leq x_{e}x_{e^{\prime}}$ by Property (2.11c).
For an edge $e\in[e^{*}]\cap\delta^{k,\mathrm{umk}}_{i}$, we have $\widehat{\operatorname*{\mathbb{E}}}\big{[}\sum_{e^{\prime}\in\delta^{k,\mathrm{mk}}_{i}}X_{e}X_{e^{\prime}}p_{e^{\prime}}\big{]}\leq\sum_{e^{\prime}\in\delta^{k,\mathrm{mk}}_{i}}x_{e}x_{e^{\prime}}p_{e^{\prime}}$ by Property (2.11e). Notice that $e^{\prime}\prec e$ for every $e^{\prime}\in\delta^{k,\mathrm{mk}}_{i}$. So, for an edge $e\in[e^{*}]\cap\delta^{k,\mathrm{umk}}_{i}$, we have
@@
The last equality follows from that we mark the first few edges of volume $\beta\rho^{k}$ in $\sigma^{k}_{i}$ in Algorithm 1 according to the order $\prec$.
Combining the inequality, (8) and that $\widehat{\operatorname*{\mathbb{E}}}\big{[}\sum_{e\preceq e^{*}}X_{e}p_{e}^{2}\big{]}=\sum_{e\preceq e^{*}}x_{e}p_{e}^{2}$, we have